OpenStudy (anonymous):

Please help. I will give medal and fan. Solve this equation : 2x-5=3(x+1)

hartnn (hartnn):

start with 3 (x+1) make use of distributive property \(a (b+c) = ab +ac\) then what is 3 (x+1) = ... ?

OpenStudy (anonymous):

3x+3

hartnn (hartnn):

correct :) so we have \(2x-5 = 3x+3 \) subtract 2x from both sides, what'd u get ?

OpenStudy (anonymous):

y do we have to subtract 2x from both sides

hartnn (hartnn):

because we combine like terms on same side :) 2x and 3x are LIKE terms to combine them, they must be on same side

hartnn (hartnn):

by subtracting 2x from both sides, you will transfer 2x from left side to right side

OpenStudy (anonymous):

okay

hartnn (hartnn):

\(2x-2x-5 = 3x-2x+3\)

OpenStudy (anonymous):

okay

hartnn (hartnn):

so what you get on both sides? 2x-2x = (2-2)x = 0

OpenStudy (anonymous):

sorry i dint understand this step

hartnn (hartnn):

a-a = 0 so 2x -2x =0

OpenStudy (anonymous):

yeah then

hartnn (hartnn):

2x−2x−5=3x−2x+3 -5 = (3-2)x + 3

OpenStudy (anonymous):

ok so whats the final ans

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

tell the final ans

OpenStudy (anonymous):

x=-2 ???

hartnn (hartnn):

(3-2)x is just x -5 = x+3 subtract 3 from both sides and you'll get your answer

hartnn (hartnn):

and its not -2

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

no 8

hartnn (hartnn):

x =-8 is correct :)

OpenStudy (anonymous):

hehehehe

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i gave u ur medal

hartnn (hartnn):

welcome ^_^